# rtboxstyle linecolor=black, backgroundcolor=lightlight-gray, leftmargin=0cm, rightmargin=0cm, skipabove=4pt, skipbelow=4pt, linewidth=0.4pt, innertopmargin=3pt, innerbottommargin=3pt, innerrightmargin=6pt, innerleftmargin=6pt, roundcorner=2

### v1

@@ 4.3 Graph Structure Generation
Existing graph generation models primarily focus on structural connectivity, often overlooking the modeling of node and edge labels [12, 71]. Many approaches also depend on domain-specific heuristics (e.g., those tailored to molecular graphs), limiting their generalizability across graph types [56]. In contrast, provenance graphs represent a new class of heterogeneous graphs, where nodes correspond to system entities such as processes, files, and network components, and edges represent relations such as read, write, clone, and send. Generating such graphs requires a domain-agnostic model that captures both structural and label information.
To this end, we adopt GraphGen [17], a domain-independent graph generation framework capable of jointly modeling node and edge labels along with structural connectivity. GraphGen converts graphs into sequences using minimum DFS codes, a canonical representation that preserves both topology and semantics. An Long Short-Term Memory (LSTM)-based model is trained to learn the joint distribution over these sequences, enabling the generation of labeled heterogeneous graphs in a domain-agnostic fashion.
However, GraphGen is typically trained on small graphs with around 100 nodes, whereas provenance graphs often exceed 10,000 nodes, making it infeasible to feed the entire graph into the model. To make the training process scalable to large graphs, we adopt a random walk sampling strategy to extract subgraphs, and then the subgraphs can be used to construct a GraphGen-compatible training dataset. During sampling, we perform a restart-based random walk. At each step, there is a probability of returning to the starting node; otherwise, the next node is uniformly selected from the current node’s neighbors. Visited nodes and their connecting edges are dynamically added to the subgraph. The walk terminates when either a predefined number of iterations is reached or the subgraph exceeds the maximum allowed number of nodes or edges. The number of walks per starting node is determined by its degree, calculated as $\texttt{factor}\times\sqrt{\texttt{degree}}$ , ensuring that high-degree nodes generate more subgraphs. After sampling, each subgraph is normalized by remapping node IDs and removing self-loops. Only subgraphs that meet the node and edge limits and remain connected are retained. The reason why we adopt restart-based random walk for subgraph sampling is its ability to balance local exploration with broader graph coverage. Unlike simple random walk, which tends to get trapped in densely connected areas, and Metropolis-Hastings random walk, which often mixes slowly and introduces sampling bias in heterogeneous graphs, restart-based random walk periodically returns to the starting node. This mechanism helps it escape local regions and sample subgraphs that better reflect the overall graph structure. The complete procedure is presented in Algorithm 1.
The training process of GraphGen involves converting graphs into minimum DFS codes, followed by learning the distribution of these sequences using a LSTM network. Given a graph dataset $\mathcal{G}=\{G_{1},G_{2},\dots,G_{m}\}$, each graph $G_{i}$ is first transformed into its minimum DFS code sequence $S_{i}=F(G_{i})=[s_{1},s_{2},\dots,s_{|E_{i}|}]$, where $|E_{i}|$ denotes the number of edges in graph $G_{i}$, and each $s_{t}$ is an edge tuple representing the $t$-th edge in the DFS traversal.
Each edge tuple $s_{t}$ is defined as:
$s_{t}=(t_{u},t_{v},L_{u},L_{e},L_{v})$
where $t_{u}$ and $t_{v}$ are the timestamps (DFS indices) of the source and target nodes, respectively; $L_{u}$ and $L_{v}$ are the labels of the source and target nodes, respectively; and $L_{e}$ is the label of the edge between the two nodes.
At each time step $t$, the LSTM updates its hidden state $h_{t}$ based on the embedding of the previous edge tuple $s_{t-1}$, and then predicts the distribution over the next edge tuple:
$h_{t}=f_{\text{trans}}(h_{t-1},f_{\text{emb}}(s_{t-1}))$
$s_{t}\sim p(s_{t}\mid h_{t})$
Here, $f_{\text{emb}}$ is the embedding function that maps edge tuples to a vector space, and $f_{\text{trans}}$ is the transition function implemented by the LSTM.
The training objective is to minimize the binary cross-entropy (BCE) loss between the predicted and ground-truth edge tuples across the entire dataset:
$\mathcal{L}=\sum_{i=1}^{m}\sum_{t=1}^{|E_{i}|}\text{BCE}(p(s_{t}\mid s_{<t}),s_{t})$
During inference, GraphGen generates new graphs using the trained LSTM model. Starting from an initial hidden state $h_{0}$ and a start-of-sequence symbol SOS, the model iteratively generates a minimum DFS code sequence $S=[s_{1},s_{2},\dots,s_{|E|}]$. The process of updating the hidden state and sampling the next edge tuple $s_{t}$ is the same as in the training phase. The generation continues until an end-of-sequence symbol EOS is produced. The resulting DFS code sequence $S$ is then converted back into a graph $G$ by reconstructing the edges based on the timestamps and labels in the sequence, yielding the generated graph structure and labels. This approach allows the model to generate graphs that reflect the structural and semantic patterns present in the training data.
4.4 Rule-based Post-processing
After the first stage, we obtain an undirected heterogeneous graph, where both nodes and edges have distinct types. However, this graph is not directly suitable as a provenance graph structure for two main reasons. First, provenance graphs are directed by nature, whereas the generated graph is undirected. Second, despite being trained on a large dataset, the heterogeneous graph generation model occasionally produces invalid connections—such as a process-read-process relation—which do not occur in real-world scenarios.
To ensure the fidelity of generated provenance graphs, we implement a rule-based post-processing pipeline. First, we convert the undirected graph into a directed graph by replacing each undirected edge $(u,v)$ with two directed edges, $(u,v)$ and $(v,u)$. We then remove edges violating domain-specific constraints, such as connections between incompatible entity types or semantically invalid interactions. Finally, we remove all isolated nodes resulting from edge removal, as such nodes do not appear in real-world provenance graphs.
The domain-specific constraints applied during post-processing vary by dataset. For example, for the Nodlink dataset, the constraints are as follows: (1) all operations must originate from process nodes; (2) file operations must terminate at file nodes; (3) process operations must terminate at process nodes; and (4) network operations must terminate at either network or file nodes. For the Cadets, Theia, and Trace datasets, the constraints are: (1) non-read operations must originate from process nodes; and (2) read operations (e.g., read, receive, load) must terminate at process nodes.
[…]
@@ 12                   Add $v$ to $G_{\text{sampled}}$ (copy attributes from $G$)
[…]
@@
21Remove self-loops from $G_{\text{sampled}}$
@@ 22 Relabel nodes to consecutive integers
25
[…]
@@
4.5 Textual Attribute Generation
After the first two stages, we obtain a directed heterogeneous graph in which nodes of different types are connected by edges of different types, and all connections conform to predefined rules. However, this graph is not yet sufficient to serve as a complete provenance graph. In real-world scenarios, nodes typically carry detailed names, such as command lines or file paths. These names are essential, as many provenance-based intrusion detection methods rely on features extracted from node names using natural language processing techniques [53, 37]. Therefore, assigning meaningful node names is crucial for supporting accurate detection.
@@ 4.5 Textual Attribute Generation
In the third stage, we leverage the natural language generation capabilities of large language models to assign names to nodes in provenance graphs. This process involves two main challenges. First, prior studies suggest that LLMs are not inherently effective at understanding graph structures  [33]. Thus, a key issue is how to transform the structure of a provenance graph into a format that is more interpretable by LLMs. Second, LLMs lack prior knowledge of typical node names in provenance graphs, and prompting them directly often results in overly simplistic outputs with limited diversity.
@@
To address the first challenge, prior work has shown that LLMs are less effective at processing graph-structured data but perform well on sequence-based inputs [69]. Motivated by this, we propose a method that converts directed heterogeneous graphs into sequences of node and edge types using depth-first search. These sequences are then input to the LLM to generate names for the corresponding nodes. Specifically, to extract complete and non-overlapping DFS paths, we use all nodes with zero in-degree as starting points for DFS traversal. However, due to the presence of self-loops in the provenance graph, some paths may not be captured using this strategy alone. To address this, we additionally include any node involved in a self-loop as a starting point. In each extracted sequence, a node is represented as a pair $(\tau_{i},n_{i})$, where $\tau_{i}\in\mathcal{T}$ denotes the node type and $n_{i}\in\mathcal{N}$ the node name. Each edge $e_{j}$ represents the interaction type between adjacent nodes. A DFS-derived sequence is defined as:
@@
To generate realistic Q&A pairs, we adopt a masking strategy that preserves structural context while requiring the model to infer node names. During the name generation process, two types of sequence are inputted to the LLM. The first type consists of sequences in which all nodes are unnamed, serving as a cold start and requiring the LLM to infer node names based solely on the order of node and edge types. The second type includes sequences where some nodes are already named, requiring the LLM to infer the remaining names using both structural order and existing node names as context. Accordingly, we define two corresponding Q&A pair formats:
@@
[…]
@@
4.6 Semantic Correctness Validation
Evaluating semantic correctness is essential for assessing the quality of synthetic provenance graphs, as it indicates whether the generated graphs accurately reflect the underlying logic and intent of real-world system behaviors. This evaluation examines whether the relationships among entities and the sequences of interactions conform to meaningful and valid operational patterns. Despite its importance, evaluating semantic correctness remains difficult due to the lack of standardized metrics and the inherent complexity in defining ground-truth semantics for diverse and dynamic system activities.
Model-based Evaluation Mechanism. To evaluate semantic correctness, we formally define a semantic unit in the graph as a triplet: the starting node name $s$, edge type $e$, and ending node name $t$. This unit is represented as $(s,e,t)$, where $s,t\in\mathcal{N}$ (the set of all node names) and $e\in\mathcal{E}$ (the set of all edge types). Concatenating these elements forms a natural language sentence, denoted $\text{Sentence}(s,e,t)$. This sentence captures the semantics of the interaction between entities, expressing that the source node $s$ performs an action of type $e$ on the target node $t$.
To assess the semantic correctness of $\text{Sentence}(s,e,t)$, we employ a BERT(Bidirectional Encoder Representations from Transformers) [13]-based encoder to obtain its embedding. This embedding captures the semantic representation of the interaction described by the triplet and serves as the basis for subsequent classification-based evaluation. Formally, let $\text{BERT}:\mathcal{S}\to\mathbb{R}^{d}$ be the encoding function that maps a sentence $\text{Sentence}(s,e,t)\in\mathcal{S}$ (the set of all possible semantic triplets in sentence form) to a dense vector $\mathbf{v}\in\mathbb{R}^{d}$. The embedding is computed as:
$\mathbf{v}=\text{BERT}(\text{Sentence}(s,e,t))$
However, embedding of $\text{Sentence}(s,e,t)$ captures only pairwise semantic relationships, resulting in a localized evaluation. To achieve graph-level assessment, we employ a multi-layer Graph Attention Network (GAT) [60]. This architecture aggregates neighborhood information through stacked self-attention layers, where each layer updates node representations by adaptively weighting features from adjacent nodes. We employ BERT to represent the text attributes of individual nodes as the initial node embeddings. Formally, for each node $v$ with text attribute $T_{v}$, its initial embedding $\mathbf{h}_{v}^{(0)}$ is derived via BERT:
$\mathbf{h}_{v}^{(0)}=\text{BERT}(T_{v})$
Let $\text{GAT}:\mathcal{V}\to\mathbb{R}^{d^{\prime}}$ denote the encoder that maps a node to its global representation in $\mathbb{R}^{d^{\prime}}$ using these initial embeddings. For a semantic unit $(s,e,t)$, we obtain the global embeddings of both the starting node $s$ and the ending node $t$:
We then concatenate the global embeddings $\mathbf{u}_{s}$ and $\mathbf{u}_{t}$ with the local embedding $\mathbf{v}$ of the semantic unit to form a comprehensive representation:
$\mathbf{z}=[\mathbf{v};\mathbf{u}_{s};\mathbf{u}_{t}]$
This comprehensive representation $\mathbf{z}$ captures both the local and global context of the semantic unit. Finally, we feed $\mathbf{z}$ into a multilayer perceptron (MLP) discriminator to determine the correctness of the semantic unit. The MLP discriminator is defined as $\text{MLP}:\mathbf{R}^{d+2d^{\prime}}\to[0,1]$. The correctness score of the semantic unit is computed as:
Correctness $\text{Correctness Score}=\text{MLP}(\mathbf{z})$
The correctness score is a value between 0 and 1, where a higher score indicates higher semantic correctness.
Contrastive Learning-based Training Framework. To train the GAT and MLP discriminator, we employ a contrastive learning approach. We use real provenance graphs as the training set, extracting all semantic units from them and labeling them as positive samples. Additionally, to provide negative samples for contrastive learning, we construct them from the positive samples using three distinct strategies to help the classifier model learn to distinguish between semantically correct and incorrect instances. The strategies are:
Subject-Object Inversion (SOI): We swap the names of the start node and the end node. This trains the model to recognize cases where the subject and object are inverted in semantics. Formally, for a semantic unit $(s,e,t)$, we generate a negative sample by inverting the subject and object:
$(s,e,t)\rightarrow(t,e,s)$
Predicate Replacement (PR): We replace the edge type with another type of edge. This trains the model to identify incorrect predicates in semantics. However, in this method, replacing the edge type with another type may still result in semantically correct cases. To guarantee the constructed semantic units are semantically incorrect, we manually classify edge types into disjoint categories. Let $\mathcal{E}=\bigcup_{i=1}^{k}\mathcal{E}_{i}$ be the partition of edge types into $k$ categories. When replacing an edge type $e\in\mathcal{E}_{i}$, we ensure the new edge type $e^{\prime}$ is from a different category, i.e., $e^{\prime}\in\mathcal{E}_{j}$ where $i\neq j$. For a semantic unit $(s,e,t)$, we generate a negative sample by replacing the predicate:
$(s,e,t)\rightarrow(s,e^{\prime},t)\quad\text{where }e^{\prime}\in\mathcal{E}_{j}\text{ and }j\neq i$
Entity Substitution (ES): We replace the name of either the start node or the end node with the name of another node. This trains the model to recognize cases where irrelevant entities appear in semantics. It is important to note that in this construction method, replacing one element of a semantic triple with another element may still result in correct semantics. This is because the replaced element and the new element may be similar. Therefore, to construct negative samples, we first cluster the node names using K-means into disjoint clusters. Let $\mathcal{V}=\bigcup_{i=1}^{m}\mathcal{V}_{i}$ be the partition of node names into $m$ clusters. When replacing a node name $s\in\mathcal{V}_{i}$ or $t\in\mathcal{V}_{i}$, we select a node $s^{\prime}\in\mathcal{V}_{j}$ or $t^{\prime}\in\mathcal{V}_{j}$ where $i\neq j$. For a semantic unit $(s,e,t)$, we generate negative samples by substituting either the subject or the object:
$(s,e,t)$ $\rightarrow(s^{\prime},e,t)\quad\text{where }s^{\prime}\in\mathcal{V}_{j}\text{ and }j\neq i$
$\text{or}\quad(s,e,t)$ $\rightarrow(s,e,t^{\prime})\quad\text{where }t^{\prime}\in\mathcal{V}_{j}\text{ and }j\neq i$
To maintain a balanced training set, we generate one negative sample for each positive sample in the real graph. Specifically, we randomly select one of the three strategies (SOI, PR, ES) to construct the negative sample. This ensures that positive and negative samples each account for 50% of the training data. For the MLP discriminator, which determines the semantic correctness of each triple, we adopt a binary cross-entropy loss. Given a batch of semantic units $\{(s_{i},e_{i},t_{i})\}_{i=1}^{N}$ with corresponding labels $y_{i}\in\{0,1\}$, the loss function is defined as:
$\mathcal{L}_{\text{MLP}}=-\frac{1}{N}\sum_{i=1}^{N}\left[y_{i}\log f_{\text{MLP}}(\mathbf{z}_{i})+(1-y_{i})\log\left(1-f_{\text{MLP}}(\mathbf{z}_{i})\right)\right]$
5 Evaluation
To thoroughly evaluate ProvSyn, we address the following research questions. Experiments on provenance graph structure generation were conducted on a machine equipped with Intel Xeon Silver CPU, 128 GB RAM, NVIDIA RTX 3090 GPUs, and Ubuntu 22.04.4 LTS. Experiments on text attribute synthesizing were performed on a machine with Intel Xeon Platinum CPU, 512 GB RAM, NVIDIA RTX A6000 GPUs, and Ubuntu 22.04.3 LTS. The research questions are:
RQ1: What is the fidelity of synthetic provenance graphs in terms of structural, textual, temporal, and embedding-based characteristics?
RQ2: What is the semantic correctness of synthetic provenance graphs?
RQ3: Can synthetic provenance graphs mitigate class imbalance in the dataset?
RQ4: When synthetic provenance graphs are used for training set augmentation in downstream intrusion detection tasks, what is the performance of the detection task?
RQ5: What are the time and resource costs of generating synthetic provenance graphs?
Further aspects of our study, including hyperparameter analysis, ablation study are discussed extensively in the Appendix C D.
5.1 Experimental Setup
Datasets. We evaluate the effectiveness of ProvSyn on two public datasets: the Nodlink Ubuntu dataset [37] and the DARPA Engagement 3 datasets [2]. The Nodlink dataset [37] we use is a simulated provenance dataset collected on Ubuntu 20.04 using Sysdig. It provides both benign logs and attack logs containing a complete multi-stage attack chain. The detailed statistics of the Nodlink provenance graphs used for training and evaluation are shown in Table I. The DARPA E3 datasets [2] are collected from an enterprise network during adversarial engagements as part of the DARPA Transparent Computing program. APT attacks are launched by a red team, while blue teams attempt to detect them through host auditing and causality analysis. TRACE, THEIA and CADETS sub-datasets are included in our evaluation. For E3 dataset, we utilize labels from ThreaTrace. As shown in Table I, the E3 dataset, compared to Nodlink, has a larger number of nodes and edges, and also exhibits greater diversity in node and edge types.
TABLE I: Comparison of dataset statistics.
Dataset Nodes Edges Node Types Edge Types
Nodlink 16,818 30,531 3 13
Cadets 89,207 276,883 6 20
Theia 77,071 202,517 3 14
Trace 74,241 113,808 7 20
Baselines. In our approach, we first use a heterogeneous graph generation model to construct the structure of the provenance graph and then employ an LLM to generate the textual attributes of the graph. In contrast, an alternative method directly calls an LLM to synthesize provenance logs and then converts these logs into a graph. This method does not require an additional heterogeneous graph generation model. We systematically evaluate and compare the provenance data synthesized by these strong LLMs with the data synthesized by ProvSyn.
To ensure a comprehensive comparison, we selected baseline models encompassing both large-parameter closed-source models and smaller-parameter open-source models. For large-parameter LLMs, we chose GPT-4o [4], Claude-3.7-Sonnet [1], and Doubao-pro-32k [3]. For smaller-parameter LLMs, we chose Qwen2.5-7B-Instruct [66] and Llama3.1-8B-Instruct [18].
To generate provenance logs using LLMs listed above, we designed structured prompts that included task descriptions, illustrative examples, and clearly defined output formats. Each dataset was paired with a customized prompt, primarily differing in the examples provided. These examples, formatted as JSON or TSV entries, contained five elements: source node type, source node name, destination node type, destination node name, and edge type. They were selected from real audit logs to ensure both representativeness and coverage of various node and edge types. For each dataset, we collected 100 responses from each LLM. After obtaining the responses, we parsed them line by line and filtered out entries that did not conform to the expected JSON or TSV format. We also discarded entries containing node or edge types outside the predefined type sets. The remaining entries were then used to construct provenance graphs for comparison with graphs generated by ProvSyn.
Configurations. For training the heterogeneous graph generation model, we sampled 20,000 subgraphs from the real provenance graphs using a random walk strategy. Among them, 16,000 subgraphs were used for training, 2,000 for validation, and 2000 for test. The model was trained for 3,000 epochs with a batch size of 32 and a learning rate of 0.003. After training, we selected the model that achieved the best performance on the validation set for use in ProvSyn. For example, on the Theia dataset, the best-performing model was obtained after 2,380 epochs and was used in the final ProvSyn pipeline.
For the text attribute synthesizing task, we used Llama3.2-3B-Instruct as the base model. We constructed a training set of 20,000 samples, consisting of 10,000 fully masked QA pairs and 10,000 partially masked QA pairs. During training, we applied the LoRA fine-tuning method, freezing most of the model parameters and updating only 0.81% of them. The model was trained for 60 steps with a learning rate of 2e-4. During inference, the temperature was set to 1.5, and the maximum number of generated tokens was set to 2048.
5.2 Fidelity Evaluation
TABLE II: Comparison of MMD metrics across different models and datasets. Lower values indicate better performance.
Dataset Model Degree Orbit Node Label Edge Label Node Label & Degree
Nodlink GPT-4o 0.75 0.94 0.11 0.75 0.89
Claude-3.7 0.56 0.85 0.07 0.92 0.77
Doubao-pro 0.58 0.82 0.18 1.50 0.59
Qwen2.5-7B 0.43 0.59 0.05 0.94 0.40
Llama3.1-8B 0.62 - 0.10 0.91 0.81
ProvSyn 0.07 0.07 0.002 0.01 0.07
Cadets GPT-4o 0.82 0.28 0.37 0.90 0.48
Claude-3.7 0.76 0.36 0.11 1.33 0.83
Doubao-pro 0.73 0.25 0.27 1.31 0.66
Qwen2.5-7B 0.76 0.39 0.67 0.77 0.33
Llama3.1-8B 0.88 0.40 0.72 0.85 0.81
ProvSyn 0.08 0.32 0.003 0.01 0.12
Theia GPT-4o 0.94 0.35 0.46 1.34 0.53
Claude-3.7 0.92 0.85 0.18 1.44 0.81
Doubao-pro 0.82 0.67 0.46 1.86 0.89
Qwen2.5-7B 0.91 0.46 0.48 1.00 0.61
Llama3.1-8B 0.68 0.87 0.53 1.01 0.66
ProvSyn 0.16 0.54 0.005 0.01 0.29
Trace GPT-4o 0.66 1.04 0.80 1.07 0.54
Claude-3.7 0.77 1.10 0.68 1.74 0.80
Doubao-pro 0.75 1.06 0.57 1.12 0.51
Qwen2.5-7B 0.24 0.23 0.44 1.06 0.47
Llama3.1-8B 0.25 0.87 1.04 1.16 0.63
ProvSyn 0.17 0.23 0.02 0.04 0.33
Structural Evaluation. To quantitatively evaluate the structural fidelity of generated graphs, we employ five Maximum Mean Discrepancy (MMD)-based metrics [19, 59] that measure distributional divergence between generated graphs and real graphs. These metrics include: (1)degree MMD for local connectivity patterns, (2)orbit MMD for higher-order structural motifs, (3)node label MMD for vertex attribute distributions, (4)edge label MMD for edge label distributions, and (5)joint node label-degree MMD for attribute-topology interactions.
To bridge the characterization of graph features with distributional divergence measurement, we leverage kernel methods to project these heterogeneous graphs into comparable representations. We construct kernel similarity matrices using two distinct Gaussian kernel formulations. The Earth Mover’s Distance-optimized Gaussian kernel is applied to categorical distribution metrics including degree MMD, node/edge label MMD, and joint node label-degree MMD, while the Euclidean-based Gaussian kernel handles continuous feature spaces in orbit MMD. The MMD² values are computed via the unbiased estimator:
$\text{MMD}^{2}=$ $\mathbb{E}_{x,x^{\prime}\sim P}[k(x,x^{\prime})]+\mathbb{E}_{y,y^{\prime}\sim Q}[k(y,y^{\prime})]$

### v2

@@ 4.1 Graph Structure Synthesis
Structure Generation Model. Existing graph generation models primarily focus on structural connectivity, often overlooking the modeling of node and edge labels [de2018molgan, you2018graph]. Many approaches also depend on domain-specific heuristics (e.g., those tailored to molecular graphs), limiting their generalizability across graph types [samanta2020nevae]. In contrast, provenance graphs represent a new class of heterogeneous graphs, where nodes correspond to system entities and edges represent system call types. Generating such graphs requires a domain-agnostic model that captures both structural and label information.
To this end, we adopt GraphGen [goyal2020graphgen], a domain-independent graph generation framework capable of jointly modeling node and edge labels along with structural connectivity. The GraphGen training process transforms each graph $G_{i}$ in a dataset $\mathcal{G}$ into its minimum DFS code sequence $S_{i}=[s_{1},s_{2},\dots,s_{|E_{i}|}]$, where each edge tuple $s_{t}=(t_{u},t_{v},L_{u},L_{e},L_{v})$ captures the DFS timestamps, node labels, and edge label. An LSTM network then learns the distribution of these sequences: at each step $t$, the hidden state $h_{t}$ is updated via the LSTM transition using the embedding of the previous edge tuple, and the model predicts the next tuple $s_{t}\sim p(s_{t}\mid h_{t})$. Training minimizes the binary cross-entropy loss $\mathcal{L}=\sum_{i}\sum_{t}\text{BCE}(p(s_{t}\mid s_{<t}),s_{t})$. During inference, the model starts from an initial state and a start symbol, then autoregressively samples a DFS code sequence until an end symbol appears. This sequence is finally converted back into a graph $G$ by reconstructing edges and labels.
Input : Graph $G=(V,E)$, starting node $u\in V$, iterations $T$, restart probability $\alpha=0.15$, node constraints $\mathit{min\_nodes},\mathit{max\_nodes}$, edge constraints $\mathit{min\_edges},\mathit{max\_edges}$
@@ 12        Add $v$ to $G_{\text{sampled}}$ (copy attributes from $G$)
14    Add edge $(\mathit{current},v)$ to $G_{\text{sampled}}$ (copy attributes from $G$)
15     $\mathit{current}\leftarrow v$
@@
21
22
23Remove self-loops from $G_{\text{sampled}}$
@@ 24 Relabel nodes to consecutive integers
27
28if $\neg\text{is\_connected}(G_{\text{sampled}})$ or $n<\mathit{min\_nodes}$ or $n>\mathit{max\_nodes}$ or $e<\mathit{min\_edges}$ or $e>\mathit{max\_edges}$ then
@@
Subgraph Sampling. However, GraphGen is typically trained on small graphs with around hunderds nodes, whereas provenance graphs contains millions of nodes, making it infeasible to feed the entire graph into the model. To make the training process scalable to large graphs, we adopt a random walk sampling strategy to extract subgraphs, and then the subgraphs can be used to construct a GraphGen-compatible training dataset. During sampling, we perform a rwr (rwr) subgraph sampling. RWR dynamically constructs subgraphs by performing random walks that probabilistically restart at the starting node, with the number of walks scaled by node degree to prioritize high-connectivity regions, and only retains connected, normalized subgraphs within predefined size constraints (Algorithm 1). Compared to other sampling methods [leskovec2006sampling, ahmed2014graph], RWR has three advantages: (1) Computational efficiency. RWR operates incrementally and does not require the full graph or global metrics, thus enabling efficient computation for provenance graphs with millions of nodes. (2) Enhanced local exploration. Compared to simple random walks, which may be confined to dense local regions, RWR incorporates restarts to return to the starting node, thereby enabling exploration across multiple neighborhoods. Unlike Metropolis–Hastings walks, which converge slowly in networks with skewed degree distributions, RWR adaptively explores local structure without requiring long mixing times [tong2006fast]. (3) Generative modeling compatibility. By sampling connected subgraphs from a seed node through restarted walks, RWR avoids disconnected fragments produced by node- or edge-based methods [leskovec2006sampling, ahmed2014graph], resulting in more coherent training instances. Moreover, the termination conditions of RWR can be defined based on subgraph size, ensuring compatibility with the size constraints of downstream generative models.
Rule-based Refinement. After structural generation, we obtain an undirected heterogeneous graph, where both nodes and edges have distinct types. However, the generated graph cannot be directly used as the structure of a provenance graph due to two main issues. First, provenance graphs are directed, whereas the generated graph is undirected. Second, despite being trained on large dataset, the heterogeneous graph generation model occasionally produces invalid connections (e.g., process-clone-network) relation, which do not occur in real-world scenarios.
To ensure the fidelity of generated provenance graphs, we design a rule-based refinement process to convert the generated graph into a valid directed provenance graph. First, we convert the undirected graph into a directed graph by replacing each undirected edge $(u,v)$ with two directed edges, $(u,v)$ and $(v,u)$. We then remove edges violating domain-specific constraints, such as connections between incompatible entity types or semantically invalid interactions. Finally, we remove all isolated nodes resulting from edge removal. The rules are dataset-specific and are derived from the structural regularities observed in the corresponding real-world provenance graphs. Specifically, we define a set of permissible event types for each ordered pair of node types based on the ground-truth graph, and any edges in the generated graphs whose event types fall outside the predefined schema are filtered out during refinement. The complete and detailed rules for Darpa E3/E5 and OpTC are presented in Appendix B (Table 5).
4.2 Textual Attribute Synthesis
After the first stage, we obtain a directed graph in which nodes of different types are connected by edges of different types, and all connections conform to predefined rules. In this stage, we focus on assigning attributes to the nodes in the generated graph to synthesize complete provenance graphs.
@@ 4.2 Textual Attribute Synthesis
We leverage the natural language generation capabilities of large language models to assign names to nodes in provenance graphs. This process involves two main challenges. First, prior studies suggest that LLMs are not inherently effective at understanding graph structures  [wang2023can, guo2023gpt4graph, jin2024large]. Thus, a key issue is how to transform the structure of a provenance graph into a format that is more interpretable by LLMs. Second, LLMs lack prior knowledge of typical node names in provenance graphs, and prompting them directly often results in overly simplistic outputs with limited diversity.
@@
To address the first challenge, prior work has shown that LLMs are less effective at processing graph-structured data but perform well on sequence-based inputs [ye2023language, zhao2023graphtext, fatemi2023talk]. Motivated by this, we propose a method that converts directed heterogeneous graphs into sequences of node and edge types using depth-first search. These sequences are then input to the LLM to generate names for the corresponding nodes. Specifically, to extract complete and non-overlapping DFS paths, we use all nodes with zero in-degree as starting points for DFS traversal. In each extracted sequence, a node is represented as a pair $(\tau_{i},n_{i})$, where $\tau_{i}\in\mathcal{T}$ denotes the node type and $n_{i}\in\mathcal{N}$ the node name. Each edge $e_{j}$ represents the interaction type between adjacent nodes. A DFS-derived sequence is defined as:
@@
To generate realistic Q&A pairs, we adopt a masking strategy that preserves structural context while requiring the model to infer node names. During the inference process, two types of sequence are inputted to the LLM. The first type consists of sequences in which all nodes are unnamed, serving as a cold start and requiring the LLM to infer node names based solely on the order of node and edge types. The second type includes sequences where some nodes are already named, requiring the LLM to infer the remaining names using both structural order and existing node names as context. Accordingly, we define two corresponding Q&A pair formats:
@@
$Q_{\text{full}}$ $=[(\tau_{1},\text{[null]})\xrightarrow{e_{1}}(\tau_{2},\text{[null]})\xrightarrow{e_{2}}\cdots\xrightarrow{e_{k-1}}(\tau_{k},\text{[null]})],$
@@
4.3 Fidelity Evaluation Framework
To comprehensively assess the fidelity of synthetic provenance graphs, we propose a framework that evaluates synthesized graphs from complementary perspectives, including structure, text attribute, temporal sequence, embedding, and semantics. Collectively, these evaluation protocols quantify the extent to which synthetic samples can serve as reliable surrogates for empirical system activities.
Structural Evaluation. To quantitatively evaluate the structural fidelity of generated graphs, we employ Maximum Mean Discrepancy (MMD)-based metrics [gretton2006kernel, thompson2022evaluation] that measure distributional divergence between generated graphs and real graphs. These metrics include: (1) degree MMD for local connectivity patterns, (2) orbit MMD for higher-order structural motifs, (3) edge label MMD for edge label distributions, and (4) node label–degree MMD for label–topology interactions.
To bridge the characterization of graph features with distributional divergence measurement, we leverage kernel methods to project these heterogeneous graphs into comparable representations. We construct kernel similarity matrices using two distinct Gaussian kernel formulations. The Earth Mover’s Distance-optimized Gaussian kernel is applied to categorical distribution metrics including degree MMD, edge label MMD, and node label-degree MMD, while the Euclidean-based Gaussian kernel handles continuous feature spaces in orbit MMD. The MMD² values are computed via the unbiased estimator, where $P$ and $Q$ denote reference and generated graph distributions, $x,x^{\prime}$ are samples from $P$, and $y,y^{\prime}$ from $Q$:
$\text{MMD}^{2}=\$$\mathbb{E}_{x,x^{\prime}\sim P}[k(x,x^{\prime})]+\mathbb{E}_{y,y^{\prime}\sim Q}[k(y,y^{\prime})]$